In a competitive market, firms sell output at a price of ₵ 20. Marginal productivity per hour of the workers is described by the equation MPL = 40 - L. What is the firm’s demand curve for Labour? If the firm can hire Labour from a competitive Labour market at a wage of ₵5 per hour, how many workers should the firm hire?


1
Expert's answer
2021-10-31T17:35:19-0400

"d_L =MRP_L = 20(40-L) =800-20L"


"800 - 20L =5"


"L = \frac{(800 - 5)}{20}"


L = "39.75" which is "40" workers